# Towards an Explainable Comparison and Alignment of Feature Embeddings

### v2

@@ 4 SPEC: A Spectral Identification of Embeddings’ Mismatches
9:  Compute eigendecomposition $\Gamma_{\psi_{1},\psi_{2}}=V\mathrm{diag}(\boldsymbol{\lambda})V^{\top}$
10:  for $i\in\{1,\ldots,n\}$ do
@@
13:  Output: Eigenvalues $\lambda_{1},\ldots,\lambda_{n}$, eigenvectors $\mathbf{u}_{1},\ldots,\mathbf{u}_{n}$.
@@
Consider normalized shift-invariant kernel $k_{1}(x,y)=\kappa_{1}(x-y)$ and $k_{2}(x^{\prime},y^{\prime})=\kappa_{2}(x^{\prime}-y^{\prime})$. Then, drawing $m$ Fourier features $\omega^{(1)}_{i}\sim\widehat{\kappa}_{1}$ and $\omega^{(2)}_{i}\sim\widehat{\kappa}_{2}$, we form the RFF-proxy kernel functions $\widehat{k}_{1},\widehat{k}_{2}$. Then, considering eigenvalues $\widehat{\lambda}_{1},\ldots,\widehat{\lambda}_{n}$ and eigenvectors $\widehat{v}_{1},\ldots,\widehat{v}_{n}$ of proxy $\widehat{\Lambda}_{\psi_{1},\psi_{2}}$, for every $\delta>0$, the following holds with probability at least $1-\delta$:
$\sum_{i=1}^{n}\bigl{\|}{\Lambda}_{\psi_{1},\psi_{2}}\widehat{v}_{i}-\widehat{\lambda}_{i}\widehat{v}_{i}\bigr{\|}^{2}_{2}\,\leq\,\frac{128\log(2/\delta)}{m}$
Proof.
We defer the proof to the Appendix A.3. ∎
As the above theorem suggests, the eigenvectors of the RFF-proxy difference kernel function $\widehat{\Lambda}_{\psi_{1},\psi_{2}}$ provide an approximation of the eigenspace of the target difference kernel function ${\Lambda}_{\psi_{1},\psi_{2}}$. On the other hand, while forming the differential covariance matrix for the proxy-RFF kernel, the dimension of $\widehat{\Gamma}_{\psi_{1},\psi_{2}}$ will be $4m\times 4m$, which is finite unlike the target shift-invariant kernel. As a result, one can apply the eigenspace equivalence in Proposition 1 to the proxy kernel function to reduce the computational complexity to $O\bigl{(}m^{3}+n\bigr{)}$ computations with $m$ features and $n$ samples.
@@ 5 SPEC-based Quantification of Embedding Differences
As discussed earlier, the eigenspace of the difference kernel matrix $\Lambda_{\psi_{1},\psi_{2}}$ provides information on the differently clustered samples by the two embeddings. Therefore, the SPEC approach motivates measuring the difference of two embeddings using the eigenspectrum of $\Lambda_{\psi_{1},\psi_{2}}$. Here, we specifically focus on the spectral radius of $\Lambda_{\psi_{1},\psi_{2}}$, i.e., its maximum absolute eigenvalue $\rho(\Lambda_{\psi_{1},\psi_{2}})=\max_{1\leq i\leq n}|\lambda_{i}(\Lambda_{\psi_{1},\psi_{2}})|$ ($\rho(A)$ denotes $A$’s spectral radius). Note that $\Lambda_{\psi_{1},\psi_{2}}$ is by definition a symmetric matrix with a zero trace, and therefore its eigenvalues are all real and add up to 0. The following definition states the difference measure, which we call SPEC-diff score:
@@
Therefore, the above proposition suggests computing the top left and right eigenvector of the $(d_{1}+d_{2})\times(d_{1}+d_{2})$ difference kernel matrix, which can be computed using the power method, and subsequently to take the gradient of the scalar function $\bigl{|}\mathbf{u}_{\text{left}}^{\top}\Gamma_{\psi_{1,\theta},\psi_{2}}\mathbf{u}_{\text{right}}\bigr{|}$ which is the absolute value of the mean of the function value for each individual sample $x_{1},\ldots,x_{n}$. This property is especially suitable for applying stochastic gradient methods.
@@ 6 Numerical Results
Experimental settings. In our experiments, we computed the SPEC differential kernel covariance matrix using $m=2000$ independent random Fourier features for a Gaussian kernel. To determine the Gaussian kernel bandwidth $\sigma$, we followed the kernel-based evaluation of generative models in (Jalali et al., 2023; Pasarkar & Dieng, 2024) and selected the embeddings bandwidths such that the difference between top eigenvalue is less than 0.01. We provide the detailed SPEC algorithm in Algorithm 1. The experiments were performed on two RTX-4090 GPUs.
Figure 3: Top 4 SPEC-identified clusters comparing CLIP and DINOv2 embeddings on 10 ImageNet classes with overlaid text labels. The last row shows the UMAP representation of the top 10 SPEC-identified clusters for each embedding.
@@
To further evaluate SPEC’s performance in comparing embeddings, we apply a typographic attack on CLIP embeddings. As studied by Materzynska et al. (2022), CLIP prioritizes text added to a custom dataset over the image content. We selected 10 classes from the ImageNet-1K dataset and overlaid different text labels directly onto the images. The top four SPEC-identified clusters are presented in Figure 3, where we observe that CLIP clusters are based on the overlaid text, whereas DINOv2 clusters them based on visual features. Additionally, the top 10 principal clusters in CLIP are not well-clustered by DINOv2, and vice versa, demonstrating that SPEC effectively highlights differences between embeddings. We compared CLIP and DINOv2 embeddings under the same settings on the ImageWoof dataset (Howard, 2019), which consists of various dog breeds from ImageNet-1K. SPEC principal clusters show that DINOv2 primarily clusters images based on dog breeds, whereas CLIP groups them based on the animals’ gestures. Additional details are provided in Figure 11 of Appendix B.3.
@@
Aligning embeddings using SPEC-align In this section, we discuss how to use the SPEC-align method to align the differential kernel covariance of two embeddings. As observed in the comparison of CLIP and DINOv2 in Figures 3 and 11, DINOv2 successfully captures certain clusters that CLIP fails to distinguish. To enhance CLIP’s performance in these tasks, we aligned CLIP with DINOv2 using the ImageNet training set. Specifically, we incorporated an alignment term into the CLIP loss function, as formulated in (6), and computed the gradient using (7). The learning parameters are detailed in the Appendix B.5.
@@ 7 Conclusion
In this paper, we proposed the spectral SPEC approach to the comparison of embedding maps. The SPEC method aims to identify groups of samples clustered by one embedding model which is not grouped by another model. We formulated a scalable algorithm with $O(n)$ computations to apply SPEC to a dataset of size $n$. We also discussed the application of SPEC for measuring the mismatches of two embeddings and their alignment. We note that the SPEC approach operates based on the assumption that the differently clustered samples can be detected by the spectral method. Extending the clustering-based approach to non-spectral clustering frameworks will be interesting for future exploration. In addition, extending the framework to compare cross-modal embeddings such as CLIP, BLIP, and ALIGN will be a future direction to this work.
@@ A.1 Proof of Theorem 1
11 1\,1 $=\,\mathrm{Tr}\Bigl{(}\frac{1}{n}K_{\psi_{2}}\Bigr{)}$
@@
[…]
Now, since $\Lambda_{\psi_{1},\psi_{2}}$ is a symmetric matrix, we can apply spectral decomposition to write it as $\Lambda_{\psi_{1},\psi_{2}}=V\mathrm{diag}(\boldsymbol{\lambda})V^{\top}$ where every row $\mathbf{v}_{i}$ of matrix $V$ is an eigenvector of $\Lambda_{\psi_{1},\psi_{2}}$ with corresponding eigenvalue $\lambda_{i}$, sorted as $\lambda_{1}\geq\cdots\geq\lambda_{n}$. Note that the eigenvalues are real and sum up to 00 as the trace of $\Lambda_{\psi_{1},\psi_{2}}$ is zero. Then, we can write:
@@
In the above, the last inequality holds as we know for every PSD matrix $A$ and vector $\mathbf{v}$, we have $\|A\mathbf{v}-\lambda\mathbf{v}\|_{2}\geq|\lambda_{j}-\lambda|\mathbf{v}$, where $\lambda_{j}$ is the eigenvalue of $A$ with the minimum absolute difference $|\lambda_{j}-\lambda|$. Therefore, the proof of Theorem 1 is complete.
@@ A.2 Proof of Proposition 1
$BA\,$ $=\,\frac{1}{n}\begin{bmatrix}\Phi_{\psi_{1}}^{\top}\vspace{2mm}\\ -\Phi_{\psi_{2}}^{\top}\end{bmatrix}\begin{bmatrix}\Phi_{\psi_{1}}&\Phi_{\psi_{2}}\end{bmatrix}$
$=\,\begin{bmatrix}\frac{1}{n}\Phi_{\psi_{1}}^{\top}\Phi_{\psi_{1}}&\frac{1}{n}\Phi_{\psi_{1}}^{\top}\Phi_{\psi_{2}}\vspace{2mm}\\ -\frac{1}{n}\Phi_{\psi_{2}}^{\top}\Phi_{\psi_{1}}&-\frac{1}{n}\Phi_{\psi_{2}}^{\top}\Phi_{\psi_{2}}\end{bmatrix}$
$=\,\begin{bmatrix}C_{\psi_{1}}&C_{\psi_{1},\psi_{2}}\vspace{2mm}\\ -C_{\psi_{1},\psi_{2}}^{\top}&-C_{\psi_{2}}\end{bmatrix}$
@@
A.3 Proof of Theorem 2

### v3

@@ 4 SPEC: A Spectral Identification of Embeddings’ Mismatches
9: Compute eigenvalues $\lambda_{1:{d_{1}+d_{2}}}$ and eigenvectors $\mathbf{v}_{1:{d_{1}+d_{2}}}$ of non-symmetric matrix $\Gamma_{\psi_{1},\psi_{2}}$
10:for $i\in\{1,\ldots,d_{1}+d_{2}\}$ do
@@
13:Output: Eigenvalues $\lambda_{1},\ldots,\lambda_{d_{1}+d_{2}}$, eigenvectors $\mathbf{u}_{1},\ldots,\mathbf{u}_{d_{1}+d_{2}}$.
@@
Consider normalized shift-invariant kernel $k_{1}(x,y)=\kappa_{1}(x-y)$ and $k_{2}(x^{\prime},y^{\prime})=\kappa_{2}(x^{\prime}-y^{\prime})$. Then, drawing $m$ Fourier features $\omega^{(1)}_{i}\sim\widehat{\kappa}_{1}$ and $\omega^{(2)}_{i}\sim\widehat{\kappa}_{2}$, we form the RFF-proxy kernel functions $\widehat{k}_{1},\widehat{k}_{2}$. Then, considering eigenvalues $\widehat{\lambda}_{1},\ldots,\widehat{\lambda}_{n}$ and eigenvectors $\widehat{\mathbf{v}}_{1},\ldots,\widehat{\mathbf{v}}_{n}$ of proxy $\widehat{\Lambda}_{\psi_{1},\psi_{2}}$, for every $\delta>0$, the following holds with probability at least $1-\delta$:
$\sum_{i=1}^{n}\Bigl{\|}{\Lambda}_{\psi_{1},\psi_{2}}\widehat{\mathbf{v}}_{i}-\widehat{\lambda}_{i}\widehat{\mathbf{v}}_{i}\Bigr{\|}^{2}_{2}\,\leq\,\frac{128\log(3/\delta)}{m}.$
Theorem 2 shows the eigenvectors of the RFF-proxy kernel difference $\widehat{\Lambda}_{\psi_{1},\psi_{2}}$ provide a proxy for the eigenspace of the target kernel difference function ${\Lambda}_{\psi_{1},\psi_{2}}$. Note that the differential covariance matrix of the proxy-RFF kernel, the dimension of $\widehat{\Gamma}_{\psi_{1},\psi_{2}}$ will be $4m\times 4m$, that is finite for every shift-invariant kernel. As a result, one can apply the eigenspace correspondence in Proposition 1 to the proxy kernel function and reduce the computational complexity to $O\bigl{(}m^{2}n+m^{3}\bigr{)}$ for $m$ RFF features and $n$ samples.
@@ 5 SPEC-based Quantification of Embedding Differences
As discussed earlier, the eigenspace of the kernel difference matrix $\Lambda_{\psi_{1},\psi_{2}}$ provides information on the differently clustered samples by the two embeddings. Therefore, the SPEC approach motivates measuring the difference of two embeddings using the eigenspectrum of $\Lambda_{\psi_{1},\psi_{2}}$. Here, we specifically focus on the spectral radius of $\Lambda_{\psi_{1},\psi_{2}}$, i.e., its maximum absolute eigenvalue $\rho(\Lambda_{\psi_{1},\psi_{2}})=\max_{1\leq i\leq n}|\lambda_{i}(\Lambda_{\psi_{1},\psi_{2}})|$ ($\rho(A)$ denotes $A$’s spectral radius). Note that $\Lambda_{\psi_{1},\psi_{2}}$ is by definition a symmetric matrix with a zero trace, and therefore its eigenvalues are all real and add up to 0. The following definition states the difference measure, which we call SPEC-diff score:
@@
Therefore, the above proposition suggests computing the top left and right eigenvector of the $(d_{1}+d_{2})\times(d_{1}+d_{2})$ kernel difference matrix, which can be computed using the power method, and subsequently to take the gradient of the scalar function $\bigl{|}\mathbf{u}_{\text{left}}^{\top}\Gamma_{\psi_{1,\theta},\psi_{2}}\mathbf{u}_{\text{right}}\bigr{|}$ which is the absolute value of the mean of the function value for each individual sample $x_{1},\ldots,x_{n}$. This property is especially suitable for applying stochastic gradient methods.
Remark 2.
Another viable difference function follows the $\ell_{2}$-norm of the eigenvalues of $\Lambda_{\psi_{1},\psi_{2}}$ (note that SPEC-diff is the $\ell_{\infty}$-norm of $\Lambda_{\psi_{1},\psi_{2}}$’s eigenvalues). Since $\Lambda_{\psi_{1},\psi_{2}}$ is a symmetric matrix with real eigenvalues, the squared-$\ell_{2}$-norm of its eigenvalues is its Frobenius norm-squared:
$\big{\|}\Lambda_{\psi_{1},\psi_{2}}\bigr{\|}^{2}_{F}=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\bigl{(}k_{\psi_{1}}(x_{i},x_{j})-k_{\psi_{2}}(x_{i},x_{j})\bigr{)}^{2}.$
An advantage of the above embedding distance function is the possibility of performing stochastic optimization, where given a mini-batch of $B$ samples, we can consider the estimation $\frac{1}{B^{2}}\sum_{i,j=1}^{B}\bigl{(}k_{\psi_{1}}(x_{i},x_{j})-k_{\psi_{2}}(x_{i},x_{j})\bigr{)}^{2}$ in the alignment process. This kernel difference Frobenius-norm-based alignment of embeddings has also been explored in the concurrent work by Gong et al. (2025).
@@ 6 Numerical Results
Experimental settings. In our experiments, we computed the SPEC differential kernel covariance matrix using $m=2000$ independent random Fourier features for a Gaussian kernel. To determine the Gaussian kernel bandwidth $\sigma$, we followed the kernel-based evaluation of generative models in (Jalali et al., 2023; Ospanov et al., 2024) and selected the embeddings bandwidths such that the difference between top eigenvalue is less than 0.01. We provide the detailed SPEC algorithm in Algorithm 1. The experiments were performed on two RTX-4090 GPUs.
Figure 3: Top 4 SPEC-identified clusters comparing CLIP and DINOv2 embeddings on 10 ImageNet classes with overlaid text labels.
@@
To further evaluate SPEC’s performance in comparing embeddings, we apply a typographic attack on CLIP embeddings. As studied by Materzynska et al. (2022), CLIP prioritizes text added to a custom dataset over the image content. We selected 10 classes from the ImageNet-1K dataset and overlaid different text labels directly onto the images. The top four SPEC-identified clusters are presented in Figure 3, where we observe that CLIP clusters are based on the overlaid text, whereas DINOv2 clusters them based on visual features. We also compared CLIP and DINOv2 embeddings under the same settings on the ImageWoof dataset (Howard, 2019), which consists of various dog breeds from ImageNet-1K. SPEC principal clusters show that DINOv2 primarily clusters images based on dog breeds, whereas CLIP groups them based on the animals’ gestures. Additional details are provided in Figure 15 of Appendix B.4.
@@
Aligning embeddings using SPEC-align In this section, we discuss how to use the SPEC-align method to align the differential kernel covariance of two embeddings. As shown in the comparison of CLIP and DINOv2 in Figures 3 and 15, DINOv2 captures certain clusters that CLIP fails to distinguish. To improve CLIP’s performance, we aligned it with DINOv2 using the ImageNet training set. Specifically, we added an alignment term to the CLIP loss, as formulated in (6), and computed the gradient using (7). Learning parameters are detailed in Appendix 22, and additional results using Remark 2 are provided in Appendix B.1.
@@ 7 Conclusion
In this paper, we proposed the spectral SPEC approach to the comparison of embedding maps. The SPEC method aims to identify groups of samples clustered by one embedding model which is not grouped by another model. We formulated a scalable algorithm with $O(n)$ computations to apply SPEC to a dataset of size $n$. We also discussed the application of SPEC for measuring the mismatches of two embeddings and their alignment. We note that the SPEC approach operates based on the assumption that the differently clustered samples can be detected by the spectral method. Extending the clustering-based approach to non-spectral clustering frameworks will be interesting for future exploration. In addition, extending the framework to compare cross-modal embeddings such as CLIP and BLIP is a future direction for this work.
@@ A.1 Proof of Theorem 1
$1\,$ $=\,\mathrm{Tr}\Bigl{(}\frac{1}{n}K_{\psi_{2}}\Bigr{)}$
@@
$\Bigl{\|}\Lambda_{\psi_{1},\psi_{2}}-\underbrace{\begin{bmatrix}\Lambda_{11}&\mathbf{0}\\ \mathbf{0}&\Lambda_{22}\end{bmatrix}}_{\Large\widetilde{\Lambda}_{\psi_{1},\psi_{2}}}\Bigr{\|}^{2}_{F}\,=\,2\bigl{\|}\Lambda_{12}\bigr{\|}^{2}_{F}\,\leq\,4(\epsilon^{2}_{1}+\epsilon_{2}).$
Now, since $\Lambda_{\psi_{1},\psi_{2}}$ is a symmetric matrix, we can apply spectral decomposition to write it as $\Lambda_{\psi_{1},\psi_{2}}=V\mathrm{diag}(\boldsymbol{\lambda})V^{\top}$ where every row $\mathbf{v}_{i}$ of matrix $V$ is an eigenvector of $\Lambda_{\psi_{1},\psi_{2}}$ with corresponding eigenvalue $\lambda_{i}$, sorted as $\lambda_{1}\geq\cdots\geq\lambda_{n}$. Note that the eigenvalues are real and sum up to 0 as the trace of $\Lambda_{\psi_{1},\psi_{2}}$ is zero. Then, we can write:
@@
In the above, the last inequality holds as we know for every PSD matrix $A$ and vector $\mathbf{v}$, we have $\|A\mathbf{v}-\lambda\mathbf{v}\|_{2}\geq|\lambda_{j}-\lambda|\ \|\mathbf{v}\|_{2}$, where $\lambda_{j}$ is the eigenvalue of $A$ with the minimum absolute difference $|\lambda_{j}-\lambda|$. Therefore, the proof of Theorem 1 is complete.
@@ A.2 Proof of Proposition 1
$BA\,$ $=\,\frac{1}{n}\begin{bmatrix}\Phi_{\psi_{1}}^{\top}\vskip 5.69054pt\\ -\Phi_{\psi_{2}}^{\top}\end{bmatrix}\begin{bmatrix}\Phi_{\psi_{1}}&\Phi_{\psi_{2}}\end{bmatrix}$
$=\,\begin{bmatrix}\frac{1}{n}\Phi_{\psi_{1}}^{\top}\Phi_{\psi_{1}}&\frac{1}{n}\Phi_{\psi_{1}}^{\top}\Phi_{\psi_{2}}\vskip 5.69054pt\\ -\frac{1}{n}\Phi_{\psi_{2}}^{\top}\Phi_{\psi_{1}}&-\frac{1}{n}\Phi_{\psi_{2}}^{\top}\Phi_{\psi_{2}}\end{bmatrix}$
$=\,\begin{bmatrix}C_{\psi_{1}}&C_{\psi_{1},\psi_{2}}\vskip 5.69054pt\\ -C_{\psi_{1},\psi_{2}}^{\top}&-C_{\psi_{2}}\end{bmatrix}$
@@
A.3 Computation of Eigenvectors and Eigenvalues of $\Gamma_{\psi_{1},\psi_{2}}$
Considering the notations in the previous proof, note that the defined differential covariance matrix $\Gamma_{\psi_{1},\psi_{2}}$ can be written as follows:
$\Gamma_{\psi_{1},\psi_{2}}\,=\,\begin{bmatrix}C_{\psi_{1}}&C_{\psi_{1},\psi_{2}}\vskip 5.69054pt\\ -C_{\psi_{1},\psi_{2}}^{\top}&-C_{\psi_{2}}\end{bmatrix}\,=\,\underbrace{\begin{bmatrix}I_{d_{1}\times d_{1}}&\mathbf{0}_{d_{1}\times d_{2}}\vskip 5.69054pt\\ \mathbf{0}_{d_{2}\times d_{1}}&-I_{d_{2}\times d_{2}}\end{bmatrix}}_{\text{Matrix}\>D}\underbrace{\begin{bmatrix}C_{\psi_{1}}&C_{\psi_{1},\psi_{2}}\vskip 5.69054pt\\ C_{\psi_{1},\psi_{2}}^{\top}&C_{\psi_{2}}\end{bmatrix}}_{\text{Matrix}\>C}$
In the above, $D$ is a diagonal matrix with $\pm 1$ diagonal entries, and $C$ represents the joint kernel covariance matrix of the concatenation of the embedding vectors $\phi_{1}(\psi_{1}(x)),\phi_{2}(\psi_{2}(x))]$ that is a PSD matrix. Since $C$ is a $(d_{1}+d_{2})\times(d_{1}+d_{2})$ symmetric PSD matrix, we can apply the Cholesky decomposition to write $C=Z^{\top}Z$ for an upper-triangular matrix $Z\in\mathbb{R}^{(d_{1}+d_{2})\times(d_{1}+d_{2})}$. Therefore, we have $\Gamma_{\psi_{1},\psi_{2}}=DZ^{\top}Z$. While $\Gamma_{\psi_{1},\psi_{2}}=DZ^{\top}Z$ is not a symmetric matrix, we note that the multiplication-order-flipped $\Theta=ZDZ^{\top}\in\mathbb{R}^{(d_{1}+d_{2})\times(d_{1}+d_{2})}$ will be a symmetric matrix. Therefore, the symmetric matrix $\Theta$ shares the same non-zero eigenvalues with $\Gamma_{\psi_{1},\psi_{2}}$, and for each eigenvector $w$ of $\Theta$, the vector $u=DZ^{\top}w$ will be the corresponding eigevector of $\Gamma_{\psi_{1},\psi_{2}}$. To conclude, one can perform the eigendecomposition for the symmetric matrix $\Theta$ and then convert the non-zero eigenvalues and eigenvectors to compute those of $\Gamma_{\psi_{1},\psi_{2}}$, resulting in the following procedure:
  1. 1.
    Form the PSD matrix $C=\begin{bmatrix}C_{\psi_{1}}&C_{\psi_{1},\psi_{2}}\vskip 5.69054pt\\ C_{\psi_{1},\psi_{2}}^{\top}&C_{\psi_{2}}\end{bmatrix}$
  2. 2.
    Apply Cholesky decomposition to write $C=Z^{\top}Z$ for an upper-triangular matrix $Z\in\mathbb{R}^{(d_{1}+d_{2})\times(d_{1}+d_{2})}$
  3. 3.
    Compute the symmetric matrix $\Theta=Z\begin{bmatrix}I_{d_{1}\times d_{1}}&\mathbf{0}_{d_{1}\times d_{2}}\vskip 5.69054pt\\ \mathbf{0}_{d_{2}\times d_{1}}&-I_{d_{2}\times d_{2}}\end{bmatrix}Z^{\top}$
  4. 4.
    Perform eigendecomposition on symmetric matrix $\Theta$ to compute $d_{1}+d_{2}$ pairs of eigenvalues and eigenvectors $(\lambda_{i},w_{i})_{i=1}^{d_{1}+d_{2}}$
  5. 5.
    For each $i\in\{1,\ldots d_{1}+d_{2}\}$, compute $u_{i}=\begin{bmatrix}I_{d_{1}\times d_{1}}&\mathbf{0}_{d_{1}\times d_{2}}\vskip 5.69054pt\\ \mathbf{0}_{d_{2}\times d_{1}}&-I_{d_{2}\times d_{2}}\end{bmatrix}Z^{\top}w_{i}$
A.4 Proof of Theorem 2